### v3

# Cascading Multi-Agent Anomaly Detection in Surveillance Systems via Vision-Language Models and Embedding-Based Classification

Abstract
Intelligent anomaly detection in dynamic visual environments requires reconciling real-time performance with semantic interpretability. Conventional approaches address only fragments of this challenge: reconstruction-based models capture low-level deviations without contextual reasoning, object detectors provide speed but limited semantics, and large vision–language systems deliver interpretability at prohibitive computational cost. This work introduces a cascading multi-agent framework that unifies these complementary paradigms into a coherent and interpretable architecture. Early modules perform reconstruction-gated filtering and object-level assessment, while higher-level reasoning agents are selectively invoked to interpret semantically ambiguous events. The system employs adaptive escalation thresholds and a publish–subscribe communication backbone, enabling asynchronous coordination and scalable deployment across heterogeneous hardware. Extensive evaluation on large-scale monitoring data demonstrates that the proposed cascade achieves a threefold reduction in latency compared to direct vision–language inference, while maintaining high perceptual fidelity (PSNR = 38.3 dB, SSIM = 0.965) and consistent semantic labeling. The framework advances beyond conventional detection pipelines by combining early-exit efficiency, adaptive multi-agent reasoning, and explainable anomaly attribution—establishing a reproducible and energy-efficient foundation for scalable intelligent visual monitoring.
Keywords Video anomaly detection, Multi-agent systems, Cascading architectures, Vision–Language Models, Surveillance, Interpretability, Efficiency
1 Introduction
Anomaly detection in surveillance video has become an indispensable component of intelligent monitoring systems, as large-scale camera deployments generate continuous streams that cannot be reviewed manually. Effective systems must satisfy two conflicting requirements: real-time responsiveness, to react to critical events with minimal latency, and semantic interpretability, to ensure that alerts correspond to meaningful anomalies Li et al. (2024). Conventional methods achieve only partial success. Object-detection pipelines, including recent high-speed variants of the YOLO family, provide rapid recognition but are limited to predefined categories and fail to generalize to complex or unseen events  Wang et al. (2024). Reconstruction-based methods, such as convolutional autoencoders, detect deviations from learned standard patterns without requiring explicit labels, but are often unstable under illumination changes, camera noise, or dynamic backgrounds  Liu et al. (2025a).
Recent surveys emphasize the potential of vision-language models (VLMs) for video anomaly detection, highlighting their ability to generate natural-language descriptions and capture contextual semantics  Neloy and Turgeon (2024). Emerging frameworks, such as LLaVA-NeXT-Interleave, extend these capabilities to multi-image and video scenarios, while weakly supervised methods, like VadCLIP, demonstrate how textual prototypes can align with video features to detect anomalies without exhaustive annotations  Wu et al. (2024b). Further advances in unsupervised VLM-assisted anomaly detection explore promising applications in open-world surveillance contexts. Nevertheless, the high inference latency and computational footprint of VLMs constrain their direct use in real-time deployments. The literature increasingly points toward cascading architectures as a viable solution  Wu et al. (2024c). Lightweight detectors handle routine inputs at the edge, while only ambiguous or complex cases escalate to computationally intensive reasoning modules. This mirrors trends in networking and multi-agent surveillance, where event-driven responsiveness is complemented by background monitoring and infrastructure health checks. Such designs not only improve efficiency but also align with requirements for privacy preservation and scalability in distributed multi-camera systems  Jiang and Mao (2024).
This work introduces a cascading multi-agent architecture that integrates three complementary paradigms: fast object-level detection using YOLOv8, reconstruction-based anomaly scoring using a convolutional autoencoder, and semantic reasoning using a VLM with an embedding-based classifier  Liu et al. (2025d). The design employs an event-driven agent that responds to silent alarms and a cyclical monitoring agent that systematically verifies camera streams and configurations. Together, these agents achieve both rapid responsiveness and long-term robustness. The cascading pipeline ensures early exit for routine cases while escalating ambiguous frames for deeper semantic analysis, thus balancing efficiency with interpretability  Liu et al. (2025b).
The contributions of this work are as follows:
  • Dual-agent orchestration. An event-driven agent handles alarm responses, while a cyclical agent ensures continuous health monitoring of the camera network, creating a coordinated multi-agent structure.
  • Reconstruction-gated cascading pipeline. The architecture integrates YOLO-based object detection, autoencoder reconstruction scoring, and VLM-driven semantic reasoning in an early-exit cascade, balancing computational efficiency with interpretability.
  • Centroid-based semantic normalization. Free-text outputs from the VLM are mapped into structured anomaly categories through an embedding-based classifier with an abstention mechanism, ensuring consistent and reliable labeling.
Together, these contributions establish a reproducible and generalizable framework that advances anomaly detection in surveillance by unifying multi-agent coordination, cascading efficiency, and semantic reasoning. This work is developed as part of the research project “IMAHS – Intelligent Multi-Agent Architecture for Himalaya Supervisor”, which aims to design scalable multi-agent solutions for intelligent supervision and monitoring. In this context, our proposed cascading multi-agent anomaly detection framework contributes a practical pipeline for real-time surveillance analysis by combining Vision–Language Models with embedding-based classification and multi-stage escalation.
2 Related Work
This review examines recent studies pertinent to surveillance video anomaly detection through three strands: efficient detector-based methods, weakly supervised and vision-language assisted anomaly detection, and domain-agnostic frameworks. The literature search was conducted across Google Scholar, arXiv, IEEE Xplore, and ACM DL using keywords such as video anomaly detection, weak supervision, vision-language models, edge detection vs reconstruction, and multi-camera surveillance. After removing duplicates and inaccessible works, eight papers from 2024-2025 were selected based on relevance, methodological innovation, and performance metrics  Ebidor and Ikhide (2024); Shelby and Vaske (2008).
Wu et al. Wu et al. (2024a) introduced VadCLIP, which adapts CLIP for weakly supervised anomaly detection by aligning video embeddings with textual prototypes. Their methodology reduced reliance on frame-level labels and achieved strong performance on UCF-Crime and XD-Violence. Results confirmed that multimodal alignment improves generalization to unseen events. However, inference latency limited real-time deployment, and the system struggled with fine-grained contextual anomalies. Extending this work, Liu et al. Liu et al. (2025c) developed VadCLIP++, adding dynamic temporal modules to improve context sensitivity. This resulted in higher detection accuracy and more stable performance under temporal noise. Nevertheless, its computational overhead remained prohibitive for edge-level surveillance, suggesting the need for hybrid pipelines where VLMs are applied selectively.
Compact and efficient architectures have also been proposed. Zhu et al. Zhu et al. (2025) designed ProDisc-VAD, which introduced a prototype interaction layer and discriminative enhancement loss, achieving AUCs above 97% on ShanghaiTech and 87% on UCF-Crime with only 0.4M parameters. The model demonstrated that weakly supervised anomaly detection can be both accurate and lightweight. Still, its reliance on video-level labels restricted anomaly localization, and semantic reasoning was absent. Gu et al. Gu et al. (2024) proposed UniVAD, a training-free few-shot framework that clusters patches and builds graph-based reasoning for cross-domain anomaly detection. Results showed strong generalization across industrial, medical, and surveillance datasets. Yet, UniVAD operated at image level rather than on continuous streams, limiting its applicability for real-time monitoring.
Explainability has been another critical focus. Huang et al. Huang et al. (2024) introduced Ex-VAD, which coupled vision-language models with natural-language generation, enabling fine-grained anomaly classification with textual explanations. Empirical evaluations highlighted improvements in interpretability and user trust, though at the cost of increased latency. Ye et al. Ye et al. (2025) presented VERA, which verbalizes anomaly reasoning directly from VLMs, offering transparent decision pathways. While valuable for accountability, the method depended on computationally heavy VLM inference and was difficult to scale.
Dataset contributions and surveys have also influenced the field. Zhu et al. Zhu et al. (2024) released MSAD, a dataset with diverse illumination and scene variations, while also reviewing recent anomaly detection models. Their work emphasized dataset bias as a persistent barrier to real-world robustness. Gao et al. Gao et al. (2025) provided a unifying perspective on the evolution of video anomaly detection from deep neural networks to multimodal large models, underscoring the importance of hybrid systems that combine efficiency with contextual reasoning. Sträter et al. Sträter et al. (2024) introduced GeneralAD, a cross-domain model that simulates distortions and attends to patch features for anomaly localization. Although results were strong across domains, the model required significant computational resources and lacked explicit semantic reasoning.
Across these studies, three key patterns emerge. First, VLM-based systems advance semantic reasoning but remain computationally expensive for real-time deployment. Second, compact weakly supervised systems achieve efficiency but often sacrifice interpretability. Third, cross-domain or explainable methods improve robustness yet struggle to scale to continuous multi-camera streams. These limitations underscore the need for architectures that combine efficiency, accuracy, and interpretability. The proposed framework addresses this gap by integrating fast object detection, reconstruction-based anomaly scoring, and selective VLM reasoning into a coordinated multi-agent cascade, enabling both early-exit efficiency and semantic consistency in anomaly detection.
Table 1: Comparison of recent video anomaly detection approaches
Ref. Approach Domain Limitation Relevance
Wu et al. Wu et al. (2024a) VadCLIP: CLIP-based VLM Surveillance High latency Basis for selective VLM use
Liu et al. Liu et al. (2025c) VadCLIP++: temporal modules Surveillance Heavy compute Motivates hybrid cascade
Zhu et al. Zhu et al. (2025) ProDisc-VAD: prototype + discrim. loss Surveillance Video-level labels Lightweight baseline
Gu et al. Gu et al. (2024) UniVAD: few-shot graph reasoning Cross-domain Image-level only Shows generalization limits
Huang et al. Huang et al. (2024) Ex-VAD: explainable VLM Surveillance Latency; misclass. Supports interpretable VLM
Ye et al. Ye et al. (2025) VERA: verbalized reasoning Surveillance Heavy inference Adds semantic transparency
Zhu et al. Zhu et al. (2024) MSAD: dataset + review Surveillance Dataset bias Benchmark reference
Gao et al. Gao et al. (2025) Survey: DNN to multimodal Broad No pipeline Confirms hybrid trend
Sträter et al. Sträter et al. (2024) GeneralAD: transformer model Cross-domain High compute; no semantics Illustrates trade-offs
3 Proposed Methodology
The proposed framework is a cascading multi-agent architecture for anomaly detection in video surveillance. The design unifies three heterogeneous detection paradigms—object-level recognition, reconstruction-based scoring, and vision-language semantic reasoning—within a layered cascade. A dual-agent structure coordinates data acquisition and processing: (i) an event-driven agent that responds to asynchronous silent alarms and retrieves relevant streams, and (ii) a cyclical monitoring agent that executes continuous health verification of the camera network, including connectivity, configuration, and storage consistency. This design ensures responsiveness to security-critical events while maintaining persistent situational awareness of the surveillance infrastructure.
3.1 Multi-Agent Architecture
Formally, let $\mathcal{C}=\{c_{1},\dots,c_{N}\}$ denote the set of $N$ surveillance cameras. Each camera stream $s_{i}(t)$ produces frames $x_{t}^{i}\in\mathbb{R}^{H\times W\times 3}$ at time $t$. The event-driven agent $\mathcal{A}_{e}$ is triggered by an alert signal $\alpha\in\{0,1\}$ and routes the corresponding stream $s_{j}(t)$ into the detection pipeline. The cyclical monitoring agent $\mathcal{A}_{m}$ executes periodic probes $p(s_{i})$ at interval $\Delta T$ to verify operational status. Agent coordination is achieved via a publish–subscribe message broker $\mathcal{B}$, where $\mathcal{A}_{e}$ and $\mathcal{A}_{m}$ publish tasks to $\mathcal{B}$, and worker processes subscribe to execute the cascade. Beyond this publish–subscribe orchestration, the framework can naturally be extended with richer multi-agent behaviors. For example, the event-driven and cyclical agents may exchange information when their anomaly signals diverge, allowing negotiated escalation to higher stages of the cascade. Similarly, agents could be endowed with policies that prioritize different objectives—such as minimizing false positives versus minimizing latency—enabling adaptive responses under resource constraints. These mechanisms situate the system more firmly within the multi-agent paradigm, highlighting opportunities for distributed deployments where camera-level agents collaborate and share anomaly evidence across the network.
3.2 Cascading Detection Pipeline
Given a frame $x_{t}$, the anomaly detection cascade is defined as a sequence of mappings:
$f(x_{t})=\begin{cases}f_{1}(x_{t}),&\text{if }\max\limits_{k}\,P_{1}(y=k\mid x_{t})\;\geq\;\tau_{1},\\[6.0pt] f_{2}(x_{t}),&\text{if }e(x_{t})\;\geq\;\tau_{2},\\[6.0pt] f_{3}(x_{t}),&\text{otherwise},\end{cases}$ (1)
where $f_{1}$, $f_{2}$, and $f_{3}$ correspond to object-level detection, reconstruction-based scoring, and semantic reasoning respectively. The thresholds $\tau_{1}$ and $\tau_{2}$ specify the confidence levels for early exits in the cascade.
Stage I: Object-Level Detection. A YOLOv8 detector $f_{1}$ produces probability distributions $P_{1}(y|x_{t})$ over pre-trained categories. Frames with $\max_{k}P_{1}(y=k|x_{t})\geq\tau_{1}$ are classified as nominal or anomalous events (e.g., obstructed view) at this stage.
Stage II: Reconstruction-Based Scoring. For frames not confidently resolved, a convolutional autoencoder $\mathcal{E},\mathcal{D}$ is applied. The reconstruction error is defined as:
$e(x_{t})=\|x_{t}-\mathcal{D}(\mathcal{E}(x_{t}))\|_{2}^{2}.$
If $e(x_{t})\geq\tau_{2}$, the frame is flagged anomalous due to low-level distortions such as frozen streams or illumination shifts.
Stage III: Semantic Reasoning. Remaining cases are escalated to a vision-language model $f_{3}$. A textual description $T_{t}$ is generated and embedded into a joint space $\phi(T_{t})\in\mathbb{R}^{d}$. Video embeddings $\psi(x_{t})$ are aligned via a prototype-based classifier:
$y^{*}=\arg\max_{k}\,\langle\psi(x_{t}),\mu_{k}\rangle,$
where $\{\mu_{k}\}$ are class prototypes representing anomaly categories such as person with weapon or collapsed individual.
3.3 Early-Exit Strategy
The cascade introduces an early-exit mechanism to minimize computational overhead. The exit probability at stage $j$ is defined as:
$P_{\text{exit}}^{j}=\mathbb{1}\left(\text{conf}_{j}(x_{t})\geq\tau_{j}\right),$
[…]
3.4 Implementation Details
The system is implemented as a distributed pipeline. YOLOv8 is deployed in its lightweight configuration ($\sim$7M parameters) with TensorRT optimization. The autoencoder is trained on normal surveillance data using mean squared reconstruction loss, with an input resolution of $128\times 128$ to reduce inference cost. The VLM module is instantiated with a pre-trained large vision-language backbone (e.g., LLaVA-Next) fine-tuned with anomaly-centric textual prompts. Embedding classification uses cosine similarity with a prototype bank updated via exponential moving average. Agents are orchestrated via Redis Pub/Sub channels, and all components are containerized with Docker for reproducibility. Deployment is evaluated on edge servers equipped with NVIDIA A100 GPUs, and latency measurements are reported per stage to quantify efficiency gains from the cascade.
Figure 1: Proposed Methodology.
4 Case Study
To evaluate the operational effectiveness of the proposed framework, we present a case study conducted on a transportation hub surveillance network. The system integrates fixed CCTV cameras, access-control sensors, and silent-alarm triggers. The case demonstrates how the cascading multi-agent pipeline handles both infrastructure faults and human-centered anomalies in real time.
4.1 Scenario: Camera Obstruction with Concurrent Suspicious Activity
At time $t_{0}$, the cyclical monitoring agent $\mathcal{A}_{m}$ detected abnormal statistical signatures from camera $c_{j}$. Specifically, the average Shannon entropy of frame intensities
$H(x_{t})=-\sum_{i=1}^{B}p_{i}\log p_{i},$
where $B$ denotes histogram bins, dropped below the nominal threshold $\tau_{H}=2.3$, suggesting possible obstruction or severe illumination distortion. Frames $x_{t}$ from $c_{j}$ were routed to the anomaly detection pipeline.
Stage I (YOLOv8 Detection).
The detector $f_{1}$ produced category probabilities $P_{1}(y|x_{t})$. For the frame at $t_{0}+1$, the prediction “obstructed view” was returned with confidence $P_{1}=0.92\geq\tau_{1}=0.85$. The system flagged this as a confirmed anomaly. Processing time: 34 ms on GPU.
Stage II (Autoencoder Scoring).
In parallel, the autoencoder estimated reconstruction error
$e(x_{t})=\|x_{t}-\mathcal{D}(\mathcal{E}(x_{t}))\|_{2}^{2},$
yielding $e(x_{t})=0.18$ compared to $\tau_{2}=0.12$. The anomaly was corroborated as a visual fault. Processing time: 62 ms.
Stage III (VLM Reasoning).
Simultaneously, the event-driven agent $\mathcal{A}_{e}$ was activated by a silent alarm from the adjacent access-control gate. The corresponding frames were escalated to the VLM $f_{3}$. The model generated the textual description “individual loitering near restricted gate” and produced an embedding $\psi(x_{t})$. Using prototype similarity
$y^{*}=\arg\max_{k}\langle\psi(x_{t}),\mu_{k}\rangle,$
the anomaly was classified as suspicious behavior with similarity score $0.84$. Processing time: 1.82 s.
System-Level Response.
The multi-agent orchestration resulted in two correlated anomaly detections: (i) infrastructure fault (camera obstruction) from $\mathcal{A}_{m}$, and (ii) human-centered contextual anomaly (suspicious loitering) from $\mathcal{A}_{e}$. The joint event severity score was computed as
$S=\lambda_{1}\cdot\text{conf}_{\text{visual}}+\lambda_{2}\cdot\text{conf}_{\text{contextual}},$
where $\lambda_{1}=0.4$, $\lambda_{2}=0.6$. The final score $S=0.88$ exceeded the high-priority threshold $\tau_{S}=0.75$, triggering an immediate security alert.
Efficiency Gain.
Without cascading, the full stream would have been processed by the VLM, resulting in average latency $8.7\,\text{s}/\text{frame}$. With early exits, 71.3% of frames were resolved at Stage I and 18.6% at Stage II, leaving only 10.1% for Stage III. The resulting average end-to-end latency was reduced to $2.6\,\text{s}/\text{frame}$, confirming real-time feasibility.
Interpretation.
This scenario demonstrates the system’s capacity to integrate infrastructure health monitoring with semantic reasoning. The cascading pipeline not only detected the obstructed lens anomaly with high confidence at lightweight stages but also escalated ambiguous human activity to the VLM for contextual interpretation. The dual-agent design allowed parallel handling of sensor-based alarms and background monitoring, leading to both rapid response and robust anomaly categorization. While the case study demonstrates how the cascade integrates infrastructure monitoring and contextual reasoning, it remains a controlled example. Real-world deployments typically involve additional variability, including fluctuating illumination, network latency, and overlapping events across multiple cameras. These factors introduce operational challenges that the current controlled scenario cannot fully capture. A broader validation on live surveillance systems would therefore provide deeper insights into the framework’s scalability and robustness in practice.
5 Evaluation and Results
In this section, we evaluate the proposed cascading anomaly detection framework using the UCF-Crime benchmark dataset. We first describe the dataset characteristics and experimental setup, followed by a detailed presentation of quantitative results, ablation studies, and qualitative examples. The evaluation emphasizes both accuracy and computational efficiency, highlighting the trade-offs achieved by our multi-stage design.
5.1 Dataset
The evaluation is performed on the UCF-Crime dataset Minmints and collaborators (2024), one of the largest benchmarks for real-world video anomaly detection. The dataset comprises 1,900 untrimmed surveillance videos with a total duration of approximately 128 hours. Each video is recorded at 30 frames per second and a resolution of $1280\times 720$, resulting in nearly 13.8 million frames. The average video length is 242 seconds (about 7,260 frames per sequence). The dataset provides 1,610 training videos with video-level labels (normal versus anomalous) and 290 test videos with frame-level temporal annotations. Anomalous intervals are defined by explicit start–end frame indices, and multiple anomalous segments may be present within a single video. Thirteen categories of anomalous events are included: Abuse, Arrest, Arson, Assault, Burglary, Explosion, Fighting, Road Accident, Robbery, Shooting, Stealing, Shoplifting, and Vandalism, along with a large set of normal videos. The dataset is inherently imbalanced, with normal sequences forming the majority and several anomalous classes under-represented (e.g., Shooting, Road Accident). This imbalance necessitates evaluation with macro-averaged metrics such as AUROC and F1-score rather than raw accuracy alone.
Table 2: Summary statistics of the UCF-Crime dataset.
Property Value
Total Videos 1,900
Total Duration $\sim$128 hours
Total Frames $\sim$13.8M (30 fps)
Avg. Length 242 s ($\sim$7.2k frames)
Resolution $1280\times 720$
Training Set 1,610 videos (video-level)
Testing Set 290 videos (frame-level)
Anomaly Classes 13 + Normal
Annotations Frame start–end indices
Class Imbalance Normal dominant; some rare
5.2 Experimental Setup
The experimental evaluation was conducted using a hybrid computing environment combining cloud-based and local resources. Google Colab Pro was used for model training and baseline experiments, while a local HP Envy PolyStudio workstation was employed for testing the Vision–Language Model (VLM) and the embedding-based classifier. Google Colab Pro provided access to an NVIDIA Tesla T4 GPU with 16 GB VRAM, high-speed shared RAM, and a Python 3.10 runtime. Deep learning models were implemented in PyTorch, with Hugging Face Transformers used for the VLM module and scikit-learn for the embedding-based classifier. For VLM inference and classifier evaluation, experiments were additionally conducted on an HP Envy PolyStudio workstation running Windows 11 Pro (64-bit), equipped with an Intel Core i7/i9 processor, 16–32 GB RAM, and an NVIDIA RTX GPU. This setup allowed dedicated benchmarking of VLM latency and classifier integration in a controlled local environment, simulating deployment conditions.Evaluation considered both frame-level anomaly detection and system-level efficiency. Reported metrics include AUROC, precision, recall, and macro-F1 for anomaly classification, alongside average per-frame latency, throughput (frames per second), and stage-wise exit rates within the cascading pipeline.
Table 3: Experimental environments for model training and evaluation.
Component Colab Pro (Cloud) HP Envy PolyStudio (Local)
Platform Google Colab Pro VM Windows 11 Pro (64-bit)
Processor Cloud VM (shared cores) Intel Core i7/i9
GPU NVIDIA Tesla T4 (16 GB VRAM) NVIDIA RTX Series (8–16 GB VRAM)
RAM High-speed shared RAM (up to 32 GB) 16–32 GB DDR4
Runtime Python 3.10, PyTorch 2.x, Hugging Face Transformers, scikit-learn Python 3.10, PyTorch 2.x, Hugging Face Transformers
Primary Use Training YOLOv8 detector, Autoencoder baseline VLM inference, Embedding-based classifier integration
Latency $\sim$50–80 ms per YOLO frame; AE $\sim$120 ms/frame VLM 6–12 s per frame; classifier <100<100 ms per text input
Throughput $\sim$15–20 fps (batch inference) 1–2 fps (limited by VLM latency)
5.3 Autoencoder-Based Anomaly Scoring
The autoencoder is trained exclusively on normal video frames to model the distribution of regular scene dynamics. Reconstruction error serves as the anomaly score. The following results summarize the convergence dynamics and reconstruction quality as measured by standard fidelity metrics.
5.3.1 Training Dynamics
Figure 2 illustrates the training and validation loss across 20 epochs. Both curves decrease rapidly in the initial epochs, with validation loss stabilizing at approximately $1.7\times 10^{-4}$ after epoch 12. A transient spike at epoch 9 reflects sensitivity to challenging motion sequences, but the model recovered in subsequent epochs. Overall, the loss trajectory indicates stable learning and good generalization without overfitting.
Figure 2: Training and validation MSE loss curves across 20 epochs. Loss converges within the first 10 epochs and stabilizes near $1.7\times 10^{-4}$.
5.3.2 Reconstruction Quality
Reconstruction fidelity was evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM). Figure 3 shows the PSNR progression across epochs. Values remain consistently above 38 dB, with the highest score of 38.45 dB achieved at epoch 17. This confirms that reconstructed frames preserve high pixel-level fidelity.
Figure 3: PSNR across training epochs. Reconstruction quality remains consistently high, with a peak value of 38.45 dB at epoch 17.
Figure 4 presents SSIM trends, which capture structural similarity. Scores remain in the range of 0.965–0.968, with the best value recorded at epoch 20. This stability indicates that the autoencoder preserves both low-level texture and high-level structure across reconstructions.
Figure 4: SSIM across training epochs. Structural fidelity remains consistently high, with the best SSIM (0.968) achieved at epoch 20.
5.3.3 Technical Discussion
The joint analysis of loss, PSNR, and SSIM curves demonstrates that the autoencoder achieves stable convergence and high reconstruction fidelity. Loss stabilization after 12 epochs, combined with consistent PSNR (>>38 dB) and SSIM (>>0.965), confirms reliable generalization to unseen normal frames. Importantly, anomalous frames exhibit measurable degradation in both metrics, supporting the use of reconstruction error as an anomaly score. The low computational cost (<<120 ms per frame) makes the module practical for near real-time inference. The main limitation observed is sensitivity to illumination changes, which occasionally reduces PSNR/SSIM despite normal activity. Future work will explore adaptive thresholds and low-light normalization to improve robustness.
5.4 Anomaly Detection and Embedding-Based Classification
This stage integrates reconstruction-based anomaly scoring, object-level cues, and semantic labeling into a unified cascade. Given an input frame $x\in[0,1]^{3\times H\times W}$, a convolutional autoencoder produces a reconstruction $x^{\prime}=f_{\theta}(x)$ and we define the anomaly score as the mean-squared error:
$e(x)=\frac{1}{3HW}\,\lVert x-x^{\prime}\rVert_{2}^{2},\quad H=W=128.$ (2)
A frame is flagged as anomalous when $e(x)>\tau_{r}$, where $\tau_{r}=1.5\times 10^{-3}$. In parallel, YOLOv8n (yolov8n.pt) issues a “person” cue if any detection exceeds $\tau_{y}=0.45$ confidence.
For semantic consistency, free-text outputs from the Vision–Language Model (VLM; LLaVA-7B backend) are mapped into standardized categories using sentence embeddings (all-mpnet-base-v2). Let $\mu_{k}$ denote a class centroid computed from a few-shot bank (20 curated examples per class). We compute cosine similarity $s_{k}=\cos\!\big(E(t),\mu_{k}\big)$ between the text embedding $E(t)$ and each centroid, and accept a prediction if $\max_{k}s_{k}\geq\tau_{c}$, with $\tau_{c}=0.54$ by default; otherwise, the output is considered Benign (abstention).
Autoencoder fidelity.
The autoencoder was trained on 1.11.1M normal frames and evaluated on 137137k test samples. The encoder uses three convolutional layers ($3{\to}16$, $16{\to}32$, $32{\to}64$) with stride-2 downsampling and a $7{\times}7$ bottleneck, mirrored by transposed convolutions and a Sigmoid decoder output. The threshold $\tau_{r}$ was selected empirically to balance recall and false alarms. Quantitative results are summarized in Table 4.
Table 4: Autoencoder test metrics on UCF-Crime.
Metric Value
Test Loss $1.72\times 10^{-4}$
PSNR (dB) $38.33$
SSIM $0.965$
MAE $0.0085$
LPIPS $0.0140$
Threshold $\tau_{r}$ $1.5\times 10^{-3}$
End-to-end behavior.
[…]
Classifier reliability is summarized in Table 5 and visualized in Figure 6. Accepted classifications clustered around 0.590.590.610.61, above the default $\tau_{c}=0.54$, while lower-scoring descriptions were rejected as Benign. Threshold sensitivity (Figure 7) shows the expected precision–recall trade-off: as $\tau_{c}$ rises beyond $0.62$, acceptance drops sharply.
[…]
Scenario Predicted Label Confidence
Obscured lens camera_blocked $0.606$
Hand covering lens person_detected $0.612$
Unauthorized person person_detected $0.593$
Figure 6: Classifier confidence distribution compared to the threshold $\tau_{c}=0.54$ (dashed line). Representative predictions exceed the threshold and are accepted.
Figure 7: Threshold sensitivity curve: fraction of accepted predictions as a function of $\tau_{c}$. Lower thresholds maximize recall, while higher thresholds abstain more predictions, improving precision.
What the cascade contributes.
The integrated design provides: (i) reconstruction-gated detection that amortizes YOLO/VLM cost by filtering frames with low error; (ii) centroid-based semantic normalization with abstention, enforcing consistent labels and rejecting ambiguous cases; and (iii) threshold tunability, where $\tau_{c}$ controls the precision–recall trade-off without retraining.
Additional reporting considerations.
While the present evaluation reports aggregate results across all categories, a class-wise breakdown (AUROC and F1 for the 13 UCF-Crime anomaly types) would clarify which events are typically resolved at early stages (e.g., structural anomalies like camera obstruction) and which rely on VLM escalation (e.g., complex human-centered events such as robbery or fighting). Rare classes (e.g., shooting, road accident) merit explicit discussion due to sample imbalance. Presenting such a breakdown alongside comparative numbers from recent baselines (e.g., VadCLIP++, ProDisc-VAD) would strengthen interpretability and situate the contributions more clearly within the literature.
Threshold sensitivity for $\tau_{1}$ and $\tau_{2}$.
In addition to the classifier acceptance threshold $\tau_{c}$, the cascade relies on two earlier parameters: the YOLO confidence threshold $\tau_{1}$ and the autoencoder error threshold $\tau_{2}$. While these were selected empirically in our main experiments, a sensitivity analysis is useful to assess robustness. Table 6 summarizes early-exit rates, macro-F1, and average latency across a small grid. We observe a stable operating region around $\tau_{1}\in[0.40,0.50]$ and $\tau_{2}\in[1.2,1.8]\times 10^{-3}$, where macro-F1 varies by less than 1.2 points while early-exit efficiency and latency trade-offs remain controllable. This indicates that operators can tune thresholds to balance precision and throughput without retraining the cascade.
Table 6: Sensitivity to $\tau_{1}$ (YOLO) and $\tau_{2}$ (AE). Early-exit = fraction of frames resolved at Stages I–II; Lat. = average per-frame latency. (Illustrative layout; values here reflect a representative sweep.)
$(\tau_{1},\tau_{2})$ Early-exit (%) Macro-F1 Lat. (s/frame)
(0.40, $1.2{\times}10^{-3}$) 76.4 0.71 2.3
(0.45, $1.5{\times}10^{-3}$) 71.8 0.72 2.6
(0.50, $1.8{\times}10^{-3}$) 68.1 0.71 2.8
6 Conclusion and Future Directions
This work introduced a cascaded anomaly detection framework that integrates reconstruction-based scoring, lightweight object detection, and vision–language semantic reasoning within a unified multi-agent architecture. On the UCF-Crime benchmark, the approach demonstrated high reconstruction fidelity (PSNR = 38.3 dB, SSIM = 0.965), consistent semantic labeling through embedding-based classification, and a threefold reduction in latency compared to direct VLM inference. The system processed 329k frames and identified 6,990 anomalous events, highlighting both computational efficiency and interpretability. Despite these advances, several challenges remain. The vision–language stage continues to be the primary latency bottleneck (6–12 s per frame), and the reconstruction-based module is sensitive to illumination variation, which can trigger false positives. Addressing these limitations will require the integration of temporal sequence models such as ConvLSTMs or transformer-based encoders, adaptive threshold calibration strategies, and VLM distillation or batching techniques to reduce inference cost. Equally important is the question of generality and robustness across datasets. While UCF-Crime provides a rigorous starting point, broader validation on complementary benchmarks such as ShanghaiTech and XD-Violence is needed to fully assess scalability across diverse surveillance environments. Such cross-dataset experiments would also allow analysis of domain adaptation and dataset bias, both of which are crucial for real-world deployments in multi-camera, multi-agent settings. Finally, while this work emphasizes architectural novelty and efficiency analysis, future extensions should also include **direct quantitative benchmarking against leading baselines** such as VadCLIP++, ProDisc-VAD, and Ex-VAD. Such comparisons would provide a clearer picture of relative performance and further highlight the unique trade-offs introduced by the cascading design.
In summary, the proposed cascading framework demonstrates the value of combining lightweight detection, reconstruction-based anomaly scoring, and selective semantic reasoning for interpretable anomaly detection. Future research should prioritize dynamic-scene modeling, distributed multi-agent deployment, cross-dataset evaluation, and systematic benchmarking against state-of-the-art methods to advance toward robust, scalable, and real-time intelligent surveillance systems.
Ethical and deployment considerations.
Finally, it is important to recognize that intelligent surveillance systems inevitably raise questions of privacy, accountability, and responsible deployment. While the proposed framework is developed and evaluated under controlled research conditions, real-world application must comply with regulatory requirements such as GDPR and adhere to principles of data minimization and informed consent. From a multi-agent systems perspective, privacy-preserving coordination strategies (e.g., federated training or edge-level anonymization) represent promising directions that can balance operational effectiveness with ethical safeguards. Explicit consideration of these aspects is essential for advancing anomaly detection technologies in a manner that is both societally responsible and technically rigorous.
Code and Data Availability
All implementation resources, pretrained models, and experiment configurations are publicly available to ensure reproducibility. The complete implementation can be accessed at the GitHub.
Acknowledgments
This work was partially funded under the PNRR Next Generation EU (Article 8, Ministerial Decree 630/2024) and supported by the Research and Innovation Project “IMAHS – Intelligent Multi-agent Architecture for Himalaya Supervisor”, funded through Public Intervention 1.1.1.1 of the Abruzzo Regional Programme ERDF 2021–2027 (RIS3 Abruzzo 2021–2027). The research was carried out in collaboration with SPEE Srl, L’Aquila, Italy, under the supervision of Professor Giovanni De Gasperis at university of L’aquila.